OpenStudy (anonymous):

Can someone please show me the graph for y= 850(x-90) and what is the domain, range and y intercept?

OpenStudy (anonymous):

looking at this this is almost in slope intercept form : y = mx + b but you have y = m (x-b) so i would distribute 80: 80 x (x) - 80 x 90 to get y = 80x - 7200

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw ^^

OpenStudy (anonymous):

just sub 850